OpenStudy (anonymous):

where does the following line cross the y axis 2y=3x-6

OpenStudy (anonymous):

At the y-axis, your x = 0. Sub into the equation.

OpenStudy (anonymous):

You solve for y so you would divide each term by 2 2y = 3x - 6 -- -- -- 2 2 2 y = 3/2x - 3

OpenStudy (anonymous):

y = mx + b b is where it crosses the y-axis. y = 3/2x - 3 so it crosses at -3 91rt is also correct... just another way of finding the answer.